# SIGMA: An Efficient Heterophilous Graph Neural Network with Fast Global Aggregation

### v4

@@ Abstract
Graph neural networks (GNNs) realize great success in graph learning but suffer from performance loss when meeting heterophily, i.e. neighboring nodes are dissimilar, due to their local and uniform aggregation. Existing attempts of heterophilous GNNs incorporate long-range or global aggregations to distinguish nodes in the graph. However, these aggregations usually require iteratively maintaining and updating full-graph information, which limits their efficiency when applying to large-scale graphs. In this paper, we propose SIGMA, an efficient global heterophilous GNN aggregation integrating the structural similarity measurement SimRank. Our theoretical analysis illustrates that SIGMA inherently captures distant global similarity even under heterophily, that conventional approaches can only achieve after iterative aggregations. Furthermore, it enjoys efficient one-time computation with a complexity only linear to the node set size $\mathcal{O}(n)$. Comprehensive evaluation demonstrates that SIGMA achieves state-of-the-art performance with superior aggregation and overall efficiency. Notably, it obtains 5$\times$ acceleration on the large-scale heterophily dataset pokec with over 30 million edges compared to the best baseline aggregation.
Index Terms:
Heterophily Graph Neural Networks, SimRank
II Introduction
@@
A series of studies have revealed the importance of solely utilizing global and structural information for addressing heterophily in graphs [12, 17, 18]Suresh et al. [18] suggested that nodes surrounded with similar graph structures are likely to share the same label. We propose to measure such global similarity in the graph by the SimRank metric, which is based on the intuition that two nodes are similar if they are connected by similar nodes [19]. This can be explained by a toy example in Fig.1(a), that the two nodes representing staffs’ websites are more likely to be considered similar and assigned the same label because of their similar neighbors, i.e., respective students. Likewise, student nodes are similar if connecting to similar projects. By this means, SIGMA is able to bypass the dissimilar nodes in the local neighbors, and establish distinct relationships for similar nodes even though they are not directly connected. The effectiveness of aggregation on a realistic dataset is shown in Fig.1(b) and  1(c). In contrast to traditional neighbor-based aggregation for example the PPR, SIGMA succeeds in discovering distant nodes of the same label as the center node. In fact, we derive a theoretical interpretation in Theorem IV.2 that SIGMA is capable of capturing global relationships without iterative calculation based on pairwise random walk accumulations and prove in Corollary IV.3 that such pairwise formation is capable of addressing heterophily issues. While representing effective global relationships, SIGMA also enjoys desirable complexity in both precomputation and aggregation. A separated precomputation stage efficiently approximates the SimRank matrix with only $\mathcal{O}(d^{2})$ time overhead, where $d=m/n$. After generated, the constant SimRank matrix enables SIGMA aggregation with $\mathcal{O}(n)$ complexity during GNN training and inference. Compared to previous aggregation schemes of $\mathcal{O}(m)$ or even higher complexities, SIGMA greatly alleviates the computational bottleneck for performing global aggregation on large-scale graphs. We summarize our main contributions as follows:
@@
  • We conduct extensive experiments to showcase the superiority of SIGMA on a range of datasets with diverse domains, scales, and heterophily properties. Generally, SIGMA achieves superior performance across 12 datasets and notably, it achieves approximately 4.3×4.3$\times$4.3 × speed-up on average compared to the best baseline, GloGNN in our overall evaluation.
III Preliminaries
@@
[…]
@@
IV SIGMA for Heterophily GNN Aggregation
@@
IV-A Interpreting SimRank for heterophily
@@
Definition IV.1 (Pairwise Random Walk [19]).
@@
where $t_{u:v}^{(2\ell)}$ is one possible tour of length $2\ell$, $t_{v:w}^{(\ell)}:\{v,...,w\}$ is the sub-tour of length $\ell$ compositing the total tour, and $p(w|v,t^{(\ell)}_{v:w})$ denotes the probability a random walk starting from node $v$ and reaching at node $w$ under the tour $t^{(\ell)}_{v:w}$.
Since random walk visits the neighbors of the ego node with equal probability, $p(w|v,t^{(\ell)}_{v:w})$ can be calculated as $p(w|v,t^{(\ell)}_{v:w})=\prod_{w_{i}\in t_{v:w}}\frac{1}{|N_{w_{i}}|}$, where $w_{i}$ is the $i$-th node in tour $t^{(\ell)}_{v:w}$. Generally, a higher probability of such walks indicates strong connectivity between the source and end nodes. Note that interpreting SimRank through random walk probabilities, such as the meeting probability of two $\sqrt{c}$-walks, is common in its efficient computation [23]. Here, we leverage SimRank with the pairwise random walk to provide greater intuition for addressing graph heterophily in the subsequent theorems. Pairwise random walk paves a way for node pair aggregation that can be calculated globally, compared to conventional GNN aggregation that is interpreted as random walks with limited hops, which is highly local [9, 24]. In fact, we then show that SimRank based aggregation can be decomposed into form of such walk probabilities and thus inherently shares the global property in the following theorem.
Theorem IV.2.
@@
[…]
@@
The implication above is by two folds. Firstly, SimRank based aggregation is global, as ${\small\textbf{S}}(u,v)$ accumulates the pairwise random probabilities of $\overleftrightarrow{P}(u,v|t^{(2\ell)})$, which counts on all possible tours connecting the node pair $(u,v)$ to assign useful weights for node features, no matter how distant the nodes $u$ and $v$ are in the graph. This shows greater global property compared to traditional local aggregation based on single path random walk probabilities. Secondly, it naturally retrieves such global relationship in an one-time manner, while the neighbor-based representation requires sufficient number layers conducting local aggregations to achieve global view. Based on the above implications, we further develop Corollary IV.3 to demonstrate that non-vanishing pairwise random scores $\overleftrightarrow{P}(u,v|t^{(2\ell)})$ will grasp homophily node pairs of $u,v$ and link such property to the overall performance more directly.
Corollary IV.3.
@@
Recall that in Theorem IV.2, the updated embedding $\widehat{\textbf{Z}}_{u}$ for the central node $u$ aggregates nodes with non-zero pairwise random meeting probabilities. Despite its global capability, each tour $t$ of length $2\ell$ is of even-hop, which is inherently linked to the homophily of the message-passed node sets. Specifically, we here consider binary classification tasks and assume the probability of a neighbor holding a different label than its central node is $p$ (representing the heterophily extent). Upon message passing from node $u$ to node $v$ through $t^{(2\ell)}$, we first calculate the probability of nodes $u$ and $v$ being homophilic denoted as $H_{p}^{\ell}$. Based on the tour $t^{(2\ell)}:\{u,u^{\prime}...v^{\prime},v\}$, we can conduct the calculated of $H_{p}^{\ell}$ by recursion:
@@
Under heterophily graph settings with $p>0.5$ such that $\Delta>0$, we’ll observe two key implications as the heterophily extent $p$ grows: (1) Larger $\ell$ results in a faster increase in the homophily probability and (2) $\Delta$ increases rapidly as $p$ grows, causing $H_{p}^{\ell}$ to approach one quickly. In other words, the probability of a target node being homophilic increases as the graph becomes more heterophilic and the heterophily extent increases. Additionally, implication (1) underscores the importance of long-range connections particularly in current tour structures, which further highlights the importance of introducing the global aggregation of S in Theorem IV.2.
@@
TABLE II: Mean &\&& standard variance of node-pair similarities.
@@
IV-B SIGMA Aggregation Workflow
As SimRank assigns high scores to node relations with topological similarities, it can be regarded as a powerful global aggregation for GNN representation, guiding the network to put higher weights on such distant but homophilous node pairs. Hence, our goal is to design the effective and efficient GNN model SIGMA that exploits the desirable capability of SimRank. Based on the implication of Theorem IV.2 that SimRank already carries global similarity information implicitly, our model removes the need of iteratively aggregating and updating embeddings by incorporating the aggregation matrix such as [15, 16]. Instead, we only rely on a constant SimRank matrix S which can be efficiently calculated in precomputation. To generate node representations from graph topology and node attributes, we deploy a simple and effective heterophilous GNN architecture, derived from LINKX [17] that respectively embeds the adjacency matrix $A$ and attribute matrix $X$ by two MLPs, i.e., $\texttt{MLP}_{A}$ and $\texttt{MLP}_{X}$, then joins by a third $\texttt{MLP}_{H}$. A tunable parameter $\delta\in[0,1]$, namely the feature factor, is employed to control the combination. The node representation matrix $H$ is:
@@ IV-B SIGMA Aggregation Workflow
Hence, we’ve concluded that the matrix $\widehat{\textbf{H}}$ holds grouping effect. Due to the continuous properties of the MLP functions, we can infer that matrix $\textbf{H}=\texttt{MLP}_{H}(\widehat{\textbf{H}})$ also shares such property. Next we consider the aggregation matrix S. According to [25], its $\epsilon$-approximation can be achieved through $T$ iterations of matrix multiplication, where for $\forall\ u,v,|\textbf{S}(u,v)-\textbf{S}^{T}(u,v)|<\epsilon$, we can get $\textbf{S}^{T}=\begin{matrix}\sum_{\ell=0}^{\lceil\log_{c}\epsilon\rceil}\;\end{matrix}c^{\ell}(1-c)(\textbf{P}^{\ell})(\textbf{P}^{\top})^{\ell}$. Then we first conclude that $\forall\ell\in[1,L],\|\textbf{P}_{u}^{\ell}-\textbf{P}_{v}^{\ell}\|_{2}\rightarrow 0$ holds by math induction. when $\ell=1$, both $\textbf{P}=\textbf{D}^{-1}\textbf{A}$ and $\textbf{P}^{\top}=\textbf{AD}^{-1}$ hold since D is the row sums of A. Suppose $\|\textbf{P}_{u}^{\ell-1}-\textbf{P}_{v}^{\ell-1}\|_{2}\rightarrow 0$, we have that for $\forall i,\|\textbf{P}^{\ell}(u,i)-\textbf{P}^{\ell}(v,i)\|_{2}$
@@
[…]
@@
IV-C Complexity Optimization and Analysis
@@
Graph $G$, decay factor $c$, error threshold $\epsilon$Approximate SimRank matrix $\widehat{\textbf{S}}$
[…]
@@
IV-D Discussions of SIGMA with Related Works
@@
V Other Related Work
@@ VI-A Experiment Setup
Parameter Settings. We calculate exact SimRank score for small datasets, while adopting approximation as described in Section IV-C with $\epsilon=0.1$ and $k\in\{16,32\}$ on large-scale graphs, which is sufficient to derive high-quality aggregation coefficients (see Section VI-D for detail). The initial coefficient $\alpha=0.5$ on all the datasets. The layer number of MLPH is set to 1 and 2 for small and large datasets. We use the same dataset train/validation/test splits as in Li et al. [16]. We conduct 5 and 10 repetitive experiments on the small and large datasets, respectively. Detailed exploration on parameters including feature factor $\delta$, learning rate $r$, dropout $p$ and weight decays are elaborated in Table VI.
VI-B Performance Comparison
@@
TABLE VII: The average learning time (s) on large-scale datasets. We separately show the break down of precomputation (Pre.) and aggregation (AGG) when applicable. The overall learning efficiency are marked with first, second and third place. The average speed-up of SIGMA over other methods are shown in Avg.$\uparrow$.
Model genius arxiv penn94 twitch snap pokec Avg.$\uparrow$
@@
LINKX - - 292.3 - - 51.2 - - 49.6 - - 302.7 - - 469.1 - - 672.3 1.73×1.73$\times$1.73 ×
GloGNN - 313.6 358.7 - 126.1 134.1 - 167.1 183.5 - 739.5 783.0 - 696.3 732.9 - 1417.8 1564.7 4.30×4.30$\times$4.30 ×
@@
VI-C Scalability and Efficiency Study
@@
Learning Time. We compare the learning time of SIMGA with LINKX and GloGNN as they are all decoupling heterophilous methods and achieve most competitive performance among baselines. We denote learning time the summation of pre-computation time and training time for a fair comparison since SIGMA needs to calculate the aggregation matrix before conducting network training. We use the same training set on each dataset and run 5-repeated experiments for 500 epochs. Note that for each method, we utilize configurations and parameters corresponding to their first-tier performance after tuning. The average learning time is reported in Table VII, along with the aggregation part. It can be seen that SIGMA costs the least learning time over all the large-scale datasets, thanks to its scalable pre-computation and top-$k$ global aggregation mechanism, which aligns with our complexity analysis in Section IV-C. Besides, the aggregation costs much less time for SIGMA due to the one-time similarity measurement calculation design, compared to GloGNN’s to-be-updated measurement. On datasets such as Penn94 and pokec, SIGMA outperforms GloGNN by around 10×10$\times$10 × and 5×5$\times$5 × faster learning time correspondingly, which greatly demonstrates the efficiency and scalablity of our aggregation method. Concerning LINKX, its classification performance is outperformed by SIGMA largely, which severely hinders its wide applications.
@@
Convergence. We study the convergence time as an indicator of model efficiency in graph learning. We compare SIGMA against leading baselines, including MixHop, GCNII, LINKX, and GloGNN, in Fig.4. The results show that SIGMA achieves favorable convergence, attaining high accuracy in a short training time. Generally, SIGMA, MixHop, and LINKX converge quickly to their highest scores due to their simple designs, with SIGMA often achieving better ultimate accuracy and maintaining it more consistently than LINKX and MixHop. Additionally, SIGMA demonstrates clear advantages over GloGNN: it converges faster and reaches higher final accuracy. For instance, its convergence acceleration over GloGNN reaches 10×10$\times$10 × on the graph Penn94. These results validate that SIGMA is both highly effective and efficient, making it well-suited for large-scale heterophilous graphs with superior scalability.

### v5

@@ Abstract
[…]
I Introduction
@@
A series of studies have revealed the importance of solely utilizing global and structural information for addressing heterophily in graphs [12, 17, 18]Suresh et al. [18] suggested that nodes surrounded with similar graph structures are likely to share the same label. We propose to measure such global similarity in the graph by the SimRank metric, which is based on the intuition that two nodes are similar if they are connected by similar nodes [19]. This can be explained by a toy example in Fig.1(a), that the two nodes representing staffs’ websites are more likely to be considered similar and assigned the same label because of their similar neighbors, i.e., respective students. Likewise, student nodes are similar if connecting to similar projects. By this means, SIGMA is able to bypass the dissimilar nodes in the local neighbors, and establish distinct relationships for similar nodes even though they are not directly connected. The effectiveness of aggregation on a realistic dataset is shown in Fig.1(b) and  1(c). In contrast to traditional neighbor-based aggregation for example the PPR, SIGMA succeeds in discovering distant nodes of the same label as the center node. In fact, we derive a theoretical interpretation in Theorem III.2 that SIGMA is capable of capturing global relationships without iterative calculation based on pairwise random walk accumulations and prove in Corollary III.3 that such pairwise formation is capable of addressing heterophily issues. While representing effective global relationships, SIGMA also enjoys desirable complexity in both precomputation and aggregation. A separated precomputation stage efficiently approximates the SimRank matrix with only $\mathcal{O}(d^{2})$ time overhead, where $d=m/n$. After generated, the constant SimRank matrix enables SIGMA aggregation with $\mathcal{O}(n)$ complexity during GNN training and inference. Compared to previous aggregation schemes of $\mathcal{O}(m)$ or even higher complexities, SIGMA greatly alleviates the computational bottleneck for performing global aggregation on large-scale graphs. We summarize our main contributions as follows:
@@
  • We conduct extensive experiments to showcase the superiority of SIGMA on a range of datasets with diverse domains, scales, and heterophily properties. Generally, SIGMA achieves superior performance across 12 datasets and notably, it achieves approximately $4.3\times$ speed-up on average compared to the best baseline, GloGNN in our overall evaluation.
II Preliminaries
@@
$\textbf{A},\textbf{X},\textbf{Y}$ The input adjacency, node feature and label matrix
$\textbf{Z},\textbf{H}$ The final updated and middle aggregated node embedding matrix
@@
III SIGMA for Heterophily GNN Aggregation
@@
III-A Interpreting SimRank for heterophily
@@
Definition III.1 (Pairwise Random Walk [19]).
@@
[…]
Theorem III.2.
@@
Denoted $\textbf{S}^{\prime}(u,v)=\sum_{\ell=1}^{\infty}c^{\ell}\cdot\overleftrightarrow{P}(u,v|t^{(2\ell)})$, it can be easily verified that ${\small\textbf{S'}}(u,v)=1$ if $u=v$, and ${\small\textbf{S}^{\prime}}(u,v)=0$ if there is no tour $t$ consisting of two separate random walk paths starting from $u$ and $v$ and ends in any same node $w$. Then, by performing one step random walk, we split ${\small\textbf{S}}^{\prime}(u,v)$ as
@@
The implication above is by two folds. Firstly, SimRank based aggregation is global, as ${\small\textbf{S}}(u,v)$ accumulates the pairwise random probabilities of $\overleftrightarrow{P}(u,v|t^{(2\ell)})$, which counts on all possible tours connecting the node pair $(u,v)$ to assign useful weights for node features, no matter how distant the nodes $u$ and $v$ are in the graph. This shows greater global property compared to traditional local aggregation based on single path random walk probabilities. Secondly, it naturally retrieves such global relationship in an one-time manner, while the neighbor-based representation requires sufficient number layers conducting local aggregations to achieve global view. Based on the above implications, we further develop Corollary III.3 to demonstrate that non-vanishing pairwise random scores $\overleftrightarrow{P}(u,v|t^{(2\ell)})$ will grasp homophily node pairs of $u,v$ and link such property to the overall performance more directly.
Corollary III.3.
@@
Recall that in Theorem III.2, the updated embedding $\widehat{\textbf{Z}}_{u}$ for the central node $u$ aggregates nodes with non-zero pairwise random meeting probabilities. Despite its global capability, each tour $t$ of length $2\ell$ is of even-hop, which is inherently linked to the homophily of the message-passed node sets. Specifically, we here consider binary classification tasks and assume the probability of a neighbor holding a different label than its central node is $p$ (representing the heterophily extent). Upon message passing from node $u$ to node $v$ through $t^{(2\ell)}$, we first calculate the probability of nodes $u$ and $v$ being homophilic denoted as $H_{p}^{\ell}$. Based on the tour $t^{(2\ell)}:\{u,u^{\prime}...v^{\prime},v\}$, we can conduct the calculated of $H_{p}^{\ell}$ by recursion:
@@
Under heterophily graph settings with $p>0.5$ such that $\Delta>0$, we’ll observe two key implications as the heterophily extent $p$ grows: (1) Larger $\ell$ results in a faster increase in the homophily probability and (2) $\Delta$ increases rapidly as $p$ grows, causing $H_{p}^{\ell}$ to approach one quickly. In other words, the probability of a target node being homophilic increases as the graph becomes more heterophilic and the heterophily extent increases. Additionally, implication (1) underscores the importance of long-range connections particularly in current tour structures, which further highlights the importance of introducing the global aggregation of S in Theorem III.2.
@@
TABLE II: Mean &\& standard variance of node-pair similarities.
@@
III-B SIGMA Aggregation Workflow
As SimRank assigns high scores to node relations with topological similarities, it can be regarded as a powerful global aggregation for GNN representation, guiding the network to put higher weights on such distant but homophilous node pairs. Hence, our goal is to design the effective and efficient GNN model SIGMA that exploits the desirable capability of SimRank. Based on the implication of Theorem III.2 that SimRank already carries global similarity information implicitly, our model removes the need of iteratively aggregating and updating embeddings by incorporating the aggregation matrix such as [15, 16]. Instead, we only rely on a constant SimRank matrix S which can be efficiently calculated in precomputation. To generate node representations from graph topology and node attributes, we deploy a simple and effective heterophilous GNN architecture, derived from LINKX [17] that respectively embeds the adjacency matrix $A$ and attribute matrix $X$ by two MLPs, i.e., $\texttt{MLP}_{A}$ and $\texttt{MLP}_{X}$, then joins by a third $\texttt{MLP}_{H}$. A tunable parameter $\delta\in[0,1]$, namely the feature factor, is employed to control the combination. The node representation matrix $H$ is:
@@ III-B SIGMA Aggregation Workflow
[…]
@@
Since both matrices $\textbf{S}*$ and H have grouping effect, we have
@@
III-C Complexity Optimization and Analysis
@@
1 Graph $G$, decay factor $c$, error threshold $\epsilon$
2 Approximate SimRank matrix $\widehat{\textbf{S}}$
3 $\widehat{\textbf{S}}\leftarrow 0,\,\textbf{R}\leftarrow\textbf{I}$
4 while max(u,v) $\textbf{R}(u,v)>(1-c)\epsilon$ do
5   $\widehat{\textbf{S}}(u,v)\leftarrow\widehat{\textbf{S}}(u,v)+\textbf{R}(u,v)$
6   for all $u^{\prime}\in N_{u},v^{\prime}\in N_{v}$ do
7    $\textbf{R}(u^{\prime},v^{\prime})\leftarrow\textbf{R}(u^{\prime},v^{\prime})+c\cdot\frac{\textbf{R}(u,v)}{|N_{u^{\prime}}|\cdot|N_{v^{\prime}}|}$
8   $\textbf{R}(u,v)\leftarrow 0$ $\widehat{\textbf{S}}(u,v)\leftarrow 0$, if $\widehat{\textbf{S}}(u,v)<\frac{1}{10}\epsilon$ // prune trivial.
9 return matrix $\widehat{\textbf{S}}$
@@ 5   $\widehat{\textbf{S}}(u,v)\leftarrow\widehat{\textbf{S}}(u,v)+\textbf{R}(u,v)$
III-D Discussions of SIGMA with Related Works
@@
IV Other Related Work
@@ V-A Experiment Setup
Parameter Settings. We calculate exact SimRank score for small datasets, while adopting approximation as described in Section III-C with $\epsilon=0.1$ and $k\in\{16,32\}$ on large-scale graphs, which is sufficient to derive high-quality aggregation coefficients (see Section V-D for detail). The initial coefficient $\alpha=0.5$ on all the datasets. The layer number of MLPH is set to 1 and 2 for small and large datasets. We use the same dataset train/validation/test splits as in Li et al. [16]. We conduct 5 and 10 repetitive experiments on the small and large datasets, respectively. Detailed exploration on parameters including feature factor $\delta$, learning rate $r$, dropout $p$ and weight decays are elaborated in Table VI.
V-B Performance Comparison
@@
[…]
@@
LINKX - - 292.3 - - 51.2 - - 49.6 - - 302.7 - - 469.1 - - 672.3 $1.73\times$
GloGNN - 313.6 358.7 - 126.1 134.1 - 167.1 183.5 - 739.5 783.0 - 696.3 732.9 - 1417.8 1564.7 $4.30\times$
@@
V-C Scalability and Efficiency Study
@@
Learning Time. We compare the learning time of SIMGA with LINKX and GloGNN as they are all decoupling heterophilous methods and achieve most competitive performance among baselines. We denote learning time the summation of pre-computation time and training time for a fair comparison since SIGMA needs to calculate the aggregation matrix before conducting network training. We use the same training set on each dataset and run 5-repeated experiments for 500 epochs. Note that for each method, we utilize configurations and parameters corresponding to their first-tier performance after tuning. The average learning time is reported in Table VII, along with the aggregation part. It can be seen that SIGMA costs the least learning time over all the large-scale datasets, thanks to its scalable pre-computation and top-$k$ global aggregation mechanism, which aligns with our complexity analysis in Section III-C. Besides, the aggregation costs much less time for SIGMA due to the one-time similarity measurement calculation design, compared to GloGNN’s to-be-updated measurement. On datasets such as Penn94 and pokec, SIGMA outperforms GloGNN by around $10\times$ and $5\times$ faster learning time correspondingly, which greatly demonstrates the efficiency and scalablity of our aggregation method. Concerning LINKX, its classification performance is outperformed by SIGMA largely, which severely hinders its wide applications.
@@
Convergence. We study the convergence time as an indicator of model efficiency in graph learning. We compare SIGMA against leading baselines, including MixHop, GCNII, LINKX, and GloGNN, in Fig.4. The results show that SIGMA achieves favorable convergence, attaining high accuracy in a short training time. Generally, SIGMA, MixHop, and LINKX converge quickly to their highest scores due to their simple designs, with SIGMA often achieving better ultimate accuracy and maintaining it more consistently than LINKX and MixHop. Additionally, SIGMA demonstrates clear advantages over GloGNN: it converges faster and reaches higher final accuracy. For instance, its convergence acceleration over GloGNN reaches $10\times$ on the graph Penn94. These results validate that SIGMA is both highly effective and efficient, making it well-suited for large-scale heterophilous graphs with superior scalability.